OpenStudy (anonymous):

ya'll ready for this? precal, parabolas, foci, directrix, vertex etc... solve: x^2=64

OpenStudy (anonymous):

Okay, did you read your text book?

OpenStudy (anonymous):

why? I have notes, does that count?

OpenStudy (anonymous):

Well, do you know what the foci, directrix, and vertex are?

OpenStudy (anonymous):

Where are you stuck? It looks like you just want us to solve it for you and explain everything.

OpenStudy (anonymous):

This equation isn't even a parabola. It's a line.

OpenStudy (anonymous):

yes, i have to graph them and stuff. Well, I would like to understand how to solve the equation and graph. And omg... you're right... *facepalm*

OpenStudy (anonymous):

i don't know where to start wio, do I square root the 64 to get x=8?

OpenStudy (anonymous):

Well, square root give us \[ |x| = 8 \]And this means \(x=8\) and \(x=-8\). We have two parallel lines.

OpenStudy (anonymous):

my teacher wants us to graph this... -_-

OpenStudy (anonymous):

wait a minute, i think should ask you a parabola instead or something else.

OpenStudy (anonymous):



OpenStudy (anonymous):

well yeah, but may i ask a different one now that i realize how simple that was.

OpenStudy (anonymous):

Wait that wasnt the question lol it was x^2=6y

OpenStudy (anonymous):

@iambatman Can you do this?

OpenStudy (anonymous):

x^2=6y?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So what do you need to know about it lol

OpenStudy (anonymous):

i have to solve the equation and graph it. its in the conic section of precalculus

OpenStudy (anonymous):

Well what can you tell me about the shape :p

OpenStudy (anonymous):

can you help me? do you know how to do it?

OpenStudy (anonymous):

Yes, I believe so, it's a parabola.

OpenStudy (anonymous):

A parabola is the set of all points at an equal distance from a fixed point and a fixed line. horizontal (y term is square) \[(y-k)^2=4p(x-h)\] vertical (x term is squared) \[(x-h)^2 = 4p(y-k)\] vertex \[(h,k)\]

OpenStudy (anonymous):

Also p = distance and direction from vertex to focus, distance and opposite direction from vertex to directrix.

OpenStudy (anonymous):

This is what you pretty much need to know, to solve this problem.

OpenStudy (anonymous):

So first thing, what kind of sketch would we get, horizontal or vertical?

OpenStudy (anonymous):

yeah i know about the two equations and everything. But by looking at this equation tell me how do i know if its vertical or horizontal?

OpenStudy (anonymous):

Look at what I wrote above each of the equation.

OpenStudy (anonymous):

ohhhhhhhhhhhhhh omg i gooooooooooooooooooooooottttt ittt!!!!!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@wio was this explanation good enough? :P

OpenStudy (anonymous):

im not done tho

OpenStudy (anonymous):

What do you need to know now?

OpenStudy (anonymous):

lol, how do you use the equation?

OpenStudy (anonymous):

awww man now what do i do...

OpenStudy (anonymous):

wio what are you doing?

OpenStudy (anonymous):

What's the vertex?

OpenStudy (anonymous):

0,0

OpenStudy (anonymous):

am i wrong?

OpenStudy (anonymous):

I think we need to use \[ x^2=4py \]

OpenStudy (anonymous):

\[ 6y=4py \implies \frac32=p \]

OpenStudy (anonymous):

This means the focus is at \((0,3/2)\)

OpenStudy (anonymous):

The directrix is the line \(y=-3/2\)

OpenStudy (anonymous):

Vertex is (0,0)

OpenStudy (anonymous):

Yes, that's all of it, right?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Should be able to graph it now :P

OpenStudy (anonymous):

I'm somehow not in the mood today. Thanks for helping out.

OpenStudy (anonymous):

No worries :P

OpenStudy (anonymous):

uh, thank you... um, you deserve a medal...

OpenStudy (anonymous):

I'm to lazy to draw the graph since I'm a bit busy right now, I'll just show you a little sample how it should look... It's not precise but it should help you out! Good luck!